OpenStudy (kawii2004):

Solve. z – 9 = 27 A. 3 B. 16 C. 18 D. 36

OpenStudy (alexandervonhumboldt2):

add 9 to both sides:

OpenStudy (alexandervonhumboldt2):

z-9+9=27+9 simplify

OpenStudy (anonymous):

27+9=

OpenStudy (yolomcswagginsggg):

z−9+9=27+9 z=36

OpenStudy (yolomcswagginsggg):

answer = 36

OpenStudy (seneca1):

D.

OpenStudy (anonymous):

z - 9 + 9 = 27 +9 z = 36, D

OpenStudy (alexandervonhumboldt2):

@Seneca1 please do not give out direct answers it is against the rules and it will not help the asker

OpenStudy (kawii2004):

okay